for the following equation, state the value of the discriminant and then describe the nature of the solutions.
-19x^2+4x+3=0
does the equation have one real solution, two real solutions, or two imaginary solutions?

Answer by mananth(16946)

-19x^2+4x+3=0
...
compare with equation ax^2 +bx +c
a=-19,b=+4, c=+3
..
discriminant = b^2-4ac
D= 4^2-(4*(-19)*(3))
=244
D>0
it has two real solutions
...
m.ananth@hotmail.ca